OpenStudy (anonymous):

Let v1 = (-6, 4) and v2 = (-3, 6). Compute the following. The scalar projection of v1 onto v2. The projection of v1 onto v2.

OpenStudy (jhannybean):

\[\large \sf comp_{\vec v_2} \vec v_1 = \frac{\vec v_1 \cdot \vec v_2}{|\vec v_2|}\]\[\large \sf proj_{\vec v_2}\vec v_1=\frac{\vec v_1 \cdot \vec v_2}{|\vec v_2|^2}\cdot \vec v_2 \]

OpenStudy (jhannybean):

from here it's just plug and play.

OpenStudy (jhannybean):

\[\sf \vec v_1 \cdot \vec v_2 = ((-6)(-3)) + ((4)(6))=~?\]\[\sf |\vec v_2| = \sqrt{(-3)^2+(6)^2}=~?\]

OpenStudy (jhannybean):

@Kiritina ?

OpenStudy (anonymous):

This would be the regular projection, not the scalar projection, right? Magnitude of v2=3√5

OpenStudy (jhannybean):

the "regular projection" = \(\sf proj_{\vec v_2}\vec v_1\) scalar projection = \(\sf comp_{\vec v_2}\vec v_1\)

OpenStudy (anonymous):

Dot product=42 so the result would be 42/(3√5)^2 *v2. Should magnitude be plugged in for v2 here?

OpenStudy (jhannybean):

magnitude = \(\sf |\vec v_2|\) vector = \(\sf \vec v_2\)

OpenStudy (anonymous):

How would you multiply \[\frac{ 42 }{ (3√5)^2 }*v2\] What would you plug in for v2? Sorry for all the questions :\

OpenStudy (jhannybean):

As I've written, \(\vec v_2\) is the vector. therefore you would plug in the vector 2. The projection of \(\vec v_2\) onto \(\vec v_1\)